Find the sum $15.01 and $27.50
Answer by Edwin McCravy(20054)

  $15.01  
   27.50
  
Add the 1 and the 0 on the right hand column. 1+0=1,
so write 1 on the right below the line:

  $15.01  
   27.50
       1

Add the 0 and the 5 in the next to last column. 0+5=5.
So write 5 just left of the 1 at the bottom:

  $15.01  
   27.50
      51

Bring down the decimal

  $15.01  
   27.50
     .51

Add the 5 and the 7. 5+7=12. That's more than 10, so write
the last digit of 12, which is 2, and carry the 1 to the 
next column:

   1
  $15.01  
   27.50
    2.51

Add the 1 to carry to the 1 and the 2 in the left column
of numbers. 1+1+2=4.  Write 4 at the bottom of the left most
column of numbers:

   1
  $15.01  
   27.50
   42.51

Now bring down the dollar mark $

   1
  $15.01  
   27.50
  $42.51

That's it.  The sum is $42.51

Edwin